### v1

# Semantic Information Management in Low-Temperature Plasma Science and Technology with VIVO

Abstract
Digital research data management is increasingly integrated across universities and research institutions, addressing the handling of research data throughout its lifecycle according to the FAIR data principles (Findable, Accessible, Interoperable, Reusable). Recent emphasis on the semantic and interlinking aspects of research data, e.g., by using ontologies and knowledge graphs further enhances findability and reusability. This work presents a framework for creating and maintaining a knowledge graph specifically for low-temperature plasma (LTP) science and technology. The framework leverages a domain-specific ontology called Plasma-O, along with the VIVO software as a platform for semantic information management in LTP research. While some research fields are already prepared to use ontologies and knowledge graphs for information management, their application in LTP research is nascent. This work aims to bridge this gap by providing a framework that not only improves research data management but also fosters community participation in building the domain-specific ontology and knowledge graph based on the published materials. The results may also support other research fields in the practical use of knowledge graphs for semantic information management.
Semantic Information System,
Research Data Management,
Domain Ontology,
Knowledge Graph Construction,
Low-Temperature Plasma Science,
keywords:
volume: 0
1 Introduction
The assimilation of digital research data management (RDM) is seeing a growing trend across universities and research institutes. This is visible from the increasing adoption rate of electronic laboratory notebooks (ELNs) and other RDM software tools [Petersen et al. (2020), Fink et al. (2023), Schröder et al. (2022), Chaerony Siffa et al. (2022)]. This trend focuses not only on the technical aspects of data handling throughout the research lifecycle but also emphasizes the importance of semantic information management in enhancing the findability, understandability, and reusability of research data. RDM practices primarily focus on the technical aspects of data handling, ensuring the availability and accessibility of research data through collection, storage, archiving, and publication on institutional or domain-specific repositories [Conrad et al. (2024), Lin et al. (2024), Rahimzadeh et al. (2023)]. However, semantic information management goes beyond organizing and storing data by adding layers of meaning and context through the use ontologies, knowledge graphs, and other semantic technologies [Schneider and Šimkus (2020), Schröder et al. (2022)]. This semantic enrichment of research data unlocks its full potential for knowledge discovery and reuse, enabling researchers to identify relationships between datasets, understand the context in which data was collected, and assess its quality and relevance for specific research questions. By making research data machine-readable and understandable, semantic information management enhances the interoperability and reusability of data, aligning with the FAIR (Findable, Accessible, Interoperable, Reusable) data principles [Wilkinson et al. (2016)].
In the field of low-temperature plasma (LTP) science and technology, often characterized by multi-disciplinary research topics and how experiments are conducted (small table-top experiments as opposed to large-scale experiments as common, e.g. in high-energy physics) [Anirudh et al. (2023)], the interlinking and explicit relationships between relevant concepts can be practical. For example, a researcher may be interested in finding datasets that are obtained from a certain diagnostic method applied to study a certain plasma source. This can be modeled by assigning a relationship between the concepts of plasma source and dataset, i.e. it can be described in an ontology, which is a formal description of concepts or knowledge and their relationships within a domain. In this example, the use of an ontology further improves the findability and comprehensibility of complex data. The work on ontologies applied in the LTP field is still in its infancy. To the best of our knowledge, there has only been a couple of academic articles on ontology development in the field of computational plasma physics [Snytnikov et al. (2020), Sapetina et al. (2020)], which is adjacent to the LTP field. Notably, the first steps in that direction have been taken within the framework of developing a metadata schema and the foundation of a general knowledge graph for applied plasma physics and plasma medicine [Sack and Hoppe (2023), Becker et al. (2023), Becker et al. (2020)].
In contrast, ontologies are already immensely prominent in pharmacology, biomedical sciences, engineering, and enterprises as a way to formalize rules and processes—provide a common understanding within the corresponding domain [Woods et al. (2023), Panzarella et al. (2023), Calvo-Cidoncha et al. (2022), Yang et al. (2023)]. Ontologies have also been applied directly in RDM workflows and pipelines. da Silva et al. [Rocha da Silva et al. (2014)] have developed the Dendro software platform for preparation of research data, on which the users have the possibility to annotate their data using a given ontology. Schröder et al. [Schröder et al. (2022)] have proposed an RDM workflow utilizing the elabFTW ELN [CARPi et al. (2017)] and the PROV ontology [Belhajjame et al. (2012)] to ensure the provenance of the generated research data. Gonçalves et al. [Gonc̨alves et al. (2017)] have developed a web tool driven by an ontology to generate web forms for the acquisition of structured data. It is worth noting that, in the age of Artificial Intelligence (AI), researchers are increasingly highlighting the crucial role of ontologies in enhancing the trustworthiness and accuracy of AI systems [Allemang and Sequeda (2024), Fernandez et al. (2023), Ronzano and Nanavati (2024)].
This work outlines a framework for constructing and maintaining a knowledge graph in the LTP field. The approach leverages a domain-specific ontology and the VIVO software platform [Conlon et al. (2019), Börner et al. (2012)] to represent and manage semantic information. VIVO is an open-source current research information system that utilize ontologies as core data structures. It is popular for managing institutional research information such as researchers, their research outputs and collaboration networks. Websites running on VIVO can be found, e.g. at https://vivo.tib.eu/fis/, https://vivo.weill.cornell.edu/, and https://vivo.brown.edu/. There, VIVO is used as a platform to discover researchers and their expertise, which can be in form of research projects and published papers. Recently, VIVO has been used in managing other kinds of linked open data (LOD) such as historical architectural prints as found in https://sah.tib.eu/. Additionally, VIVO offers a user-friendly way to ingest data into the knowledge graph, allowing users without technical expertise in graph data to participate in its construction. This fosters a community-maintained, domain-specific knowledge graph.
The rest of the paper is structured as follows. In Section 2, we briefly introduce the domain-specific ontology used in this work, the VIVO software platform, and ways to integrate the domain-specific ontology and data instances into VIVO. Subsequently, we present different potential use cases of the suggested research in Section 3, specifically in the LTP field. Finally in Section 4, we conclude and discuss the outlook of the present work.
2 Methods
This section details the development and integration of a domain-specific ontology tailored for the field of low-temperature plasma (LTP) science and technology. The overarching objective is to establish a robust semantic framework that facilitates knowledge creation, discovery, and management within the LTP domain. This is achieved through two primary approaches: the construction of a specialized ontology, Plasma-O, and its subsequent integration into the VIVO software to create and maintain the knowledge graph.
2.1 Ontology for Low-Temperature Plasma Science and Technology
The domain-specific ontology presented in this work is built upon the earlier draft [Sack and Hoppe (2023), Becker et al. (2023), Becker et al. (2020)] and the concepts introduced in the existing plasma metadata schema, named Plasma-MDS, introduced by Franke et al. [Franke et al. (2020)]. There, the introduced metadata schema is intended to support the adoption of the FAIR data principles and RDM in the LTP field, specifically for the publication of datasets generated from LTP experiments. Plasma-MDS is used as the metadata standard for LTP-specific data repositories such as INPTDAT [^1] https://www.inptdat.de/ [Becker et al. (2019)] and RDPCIDAT [^2] https://rdpcidat.rub.de/. The developed ontology, however, is designed to have a broader application range, e.g. as a basis of metadata schemas for various RDM workflows and semantic knowledge creation and discovery. In this work, we focus on the part of the domain-specific ontology for the latter case and consider the first case and the complete ontology for future work. From here on, we shall refer to this domain-specific ontology as the Plasma Ontology or Plasma-O.
In contrast with Plasma-MDS that introduces Dataset as a central class or concept, Plasma-O is built around the central class PlasmaStudy to which other classes are connected. It is formally defined (not a full definition) as
$PlasmaStudy$ $\sqsubseteq\geq 1hasResearchTopic.ResearchTopic$
$\sqsubseteq\geq 1isConductedBy.\text{foaf:}Agent$
$\sqsubseteq\exists hasOutput.\left(\text{bibo:}Document\sqcup\text{vivo:}Dataset\right)$
$\sqsubseteq\exists usesDevice.DiagnosticDevice$
$\sqsubseteq\exists usesMethod.DiagnosticMethod$
$\sqsubseteq\exists hasPlasmaSource.PlasmaSource$
$\sqsubseteq\exists hasMedium.Medium$
$\sqsubseteq\exists hasTarget.Target$
This design choice is to ensure the applicability of the ontology outside the dataset publication use case, and establishes Dataset as an output of PlasmaStudy as per the designated relationship.
Another core class in Plasma-O is Device, a generic class encompassing both PlasmaSource and DiagnosticDevice. The key distinction lies in the specific relationships PlasmaSource has with other classes connected to PlasmaStudy. For instance, DiagnosticDevice lacks direct relationships with Target and Medium, which are exclusive to PlasmaSource. It is important to note that Plasma-O, by design, includes inverse properties for most object properties. This means that if PlasmaStudy connects to PlasmaSource through the hasSource property, the inverse relationship can be expressed using the isSourceOf property.
Figure 1: Selected Plasma-O classes and their relationships with each other and classes from external ontologies.
Plasma-O can be attached to other ontologies by means of reusing existing classes from them. Here, we attach Plasma-O to the VIVO ontology [Corson-Rikert et al. (2012)], which forms the backbone of the VIVO software. The VIVO ontology itself reuses classes and properties from other ontologies such as FOAF [Brickley and Miller (2014)], BIBO [D’Arcus and Giasson (2009)], and OBO [Smith et al. (2007)], to name a few. Together, these ontologies are used to describe various different concepts that can be grouped into people, activities, courses, events, organizations, equipment, research, and locations. This encompasses the general concepts in most scientific disciplines and highlights the interoperability of the different concepts from these ontologies. Figure 1 illustrates an example of selected classes from Plasma-O attached to the VIVO ontology (which includes the classes from OBO, FOAF, and BIBO ontologies) and their basic relations. Note that classes and properties that are not preceded by a namespace belong to Plasma-O.
Plasma-O, despite being able to describe general relevant concepts in LTP, is still being developed further. The presented work strives to provide public access to the ontology and related knowledge graph going—with respect to accessibility for humans and machines—far beyond storing the source files in a public repository. The suggested approach of using VIVO as a semantic knowledge management system for domain-specific data and metadata (beyond its original application scenario) is seen as transferable to other domains. With this, it potentially fosters the adoption of the FAIR data principles in different interdisciplinary research communities. The version of Plasma-O (version 0.5.0) used in the present work is available at the Plasma-MDS GitHub page [^3] https://github.com/plasma-mds/plasma-ontology.
2.2 LTP Knowledge Graph Development with VIVO
The core technology of VIVO is the ontologies. VIVO provides a user-friendly user interface (UI) that enables end users (e.g. data curators) to create and manage semantic information in the form of a knowledge graph based on the available ontologies. As mentioned in Section 2.1, VIVO comes with the VIVO ontology and various other ontologies readily facilitating the creation of a knowledge graph representing researchers, their interests, projects, and outputs. Additionally, incorporating new ontologies and/or knowledge graphs into VIVO is simplified through its add/remove RDF data feature. Accordingly, we introduce Plasma-O into VIVO and set the namespace (or prefix) to “plasma-o” for quick identification of the ownership of the classes. Note that, we use VIVO version 1.13 for this work. For a more detailed documentation to install and set up VIVO version 1.13, we refer the readers to the relevant VIVO documentation [^4] https://wiki.lyrasis.org/display/VIVODOC113x.
Figure 2: Example workflow for creating an instance of PlasmaSource.
There are several ways to populate the ontologies with instances or data on VIVO to create a knowledge graph ab intio. The most straightforward way is to use the data input feature. As a first step, we create a class group, called “plasma-o”, and pick relevant Plasma-O classes (and also relevant classes from other related ontologies) to be included in the plasma-o group. This is done to ease the data ingest process on VIVO. We can then choose the starting class we want to populate with instances. Figure 2 shows an example workflow for creating an instance of PlasmaSource. It starts from selecting the PlasmaSource class (available under the plasma-o group). Then we will be prompted to provide a name for this instance, in this case, we choose the technical name for this plasma source, which is “kINPen-sci”. Upon creating the instance, we will be directed to the newly created kINPen-sci instance page as shown in step 3 in Figure 2. Since PlasmaSource is connected to different other classes in the ontology through the corresponding properties, these connections will be automatically displayed on the kINPen-sci instance page. Here, we can proceed to link the kINPen-sci instance with either existing instances or new instances. In the example, we link the kINPen-sci instance with an already existing PlasmaStudy instance, through the isSourceOf property (shown by its label “is plasma source of” in the figure). A similar process can be done for other classes and properties. In addition to creating the instance one by one, which can be time consuming, VIVO also allows for creation of instances and linking in bulk using CSV files. This particular approach streamlines the process by allowing users to provide instance details like id, name, uri, and other relevant properties in a structured format.
Regardless of the creation method, utilizing ontologies within VIVO maximizes the reuse of existing instances to describe various concepts. This approach is applicable to all classes across all ontologies available in VIVO, and the created instances inherit the relationships defined within those ontologies. These interconnected instances form the foundation of the knowledge graph. Here, we constructed the plasma knowledge graph (Plasma-KG) by applying these workflows and incorporating data from the INPTDAT data repository [^5] https://www.inptdat.de [Becker et al. (2019)] and various publications in the LTP field.
3 Results
The primary and most general use case of VIVO is to create profiles for researchers and their research outputs, typically in the form of published academic articles and other contributions. This directly relates to the expertise these researchers possess. For instance, a researcher might be linked to several publications and listed as experienced in specific plasma sources, devices, or diagnostic methods. Beyond this, we demonstrate other potential use cases of VIVO within the LTP field, which can be adapted by other domains as well. Some of these potential use cases are demonstrated in this section.
3.1 Semantic Cataloging and Browsing of Domain-Specific Information and Data
This use case extends the general use case of VIVO that is researcher profiling. Using Plasma-O, we can create a semantic catalog for research objects relevant in the LTP field such as plasma sources and datasets generated within the framework of plasma studies. This use case attempts to semantically enrich the existing solution offered by the INPTDAT data repository [Becker et al. (2019)], where researchers can share datasets from their experiments or studies as well as the details of the plasma sources they use or develop or related applications and patents.
INPTDAT is a DKAN [^6] https://demo.getdkan.org/-based data portal utilizing the relational database technology which does not directly support relationships between data points. In contrast, ontology-based software like VIVO automatically interlinks instances of classes like PlasmaSource and Dataset based on the applied ontology, which enriches the connections between these classes and other relevant classes. This results in enhanced navigability while browsing these research objects, which is highly relevant in terms of the FAIR data principles. Figure 3 showcases an example device catalog within VIVO along with its faceted browsing feature, where researchers can browse devices used in LTP research, including plasma sources and diagnostic devices. Additionally, researchers can navigate through various lists of instances based on their class e.g. plasma studies, targets, media, etc.
Figure 3: Example of a device catalog within the plasma knowledge graph showcasing plasma sources and diagnostic devices. Faceted browsing feature enhances navigation across the semantic information landscape of the LTP domain.
3.2 Semantic Information Retrieval Using SPARQL
VIVO offers a user friendly way to ingest data into the knowledge graph. This allows users that are not experienced in the technicalities of graph data to participate in building the knowledge graph, which results in a community-maintained (domain specific) knowledge graph. Once the knowledge graph contains enough data, we can construct complex competency questions that can be implemented using SPARQL [^7] https://www.w3.org/TR/sparql11-query/. This use case is not specific to VIVO, nonetheless, VIVO offers a SPARQL query API that allows a self-contained solution for an ontology-based information management system.
While ingesting data using VIVO is straightforward for general users, formulating complex competency questions using SPARQL requires some knowledge in the query language and the ontology itself. Automated processes to represents the questions in a natural language into SPARQL queries [^8] https://github.com/LiberAI/NSpM [Yin et al. (2021)], by which non-technical users can understand, are a topic on its own and beyond the scope of the present work. Nevertheless, we demonstrate this use case in this work by constructing two competency questions that are relevant in the LTP field as follows:
  1. 1.
    “Who is experienced in PlasmaSource at Institute?”
  2. 2.
    “Which Dataset is obtained with DiagnosticMethod applied to diagnose PlasmaSource?”
The SPARQL query for the first competency question is shown in Listing 1. This query selects PlasmaSource and Institute instances with the labels “kINPen-sci” and “INP Greifswald”, respectively. It then infers instances of foaf:Person (a subclass of foaf:Agent) related to the identified PlasmaSource and Institute. The query results, presented in Table 1, reveal four foaf:Person instances (with their ids and labels) that have experience with “kINPen-sci” (instance id: n6807) and affiliated with “INP Greifswald” (instance id: n3601).
Listing 1: SPARQL query for “Who is experienced in PlasmaSource at Institute?”
PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>
PREFIX plasma-o: <http://plasma-mds.org/ontology/>
PREFIX foaf: <http://xmlns.com/foaf/0.1/>
# Who is experienced in the plasma source "kINPen-sci"
# at the institute "INP Greifswald?"
SELECT ?Person ?PersonLabel ?PlasmaSource ?Institute
WHERE
# Define the type (the class) of Person
?Person rdf:type foaf:Person .
# Get the instance with label "INP Greifswald"
?Institute rdfs:label "INP Greifswald"@en-US .
# Get the instance with label "kINPen-sci"
?PlasmaSource rdfs:label "kINPen-sci"@en-US .
# Main query
?Person plasma-o:isAffiliatedWith ?Institute .
?Person plasma-o:isExperiencedInDevice ?PlasmaSource .
# Showing the labels for Person
?Person rdfs:label ?PersonLabel .
# Returns maximum of 5 results
LIMIT 5
Table 1: SPARQL results for “Who is experienced in PlasmaSource at Institute?” (Listing 1). Note that the names are redacted for privacy reasons.
Person PersonLabel PlasmaSource Institute
n1951 N***, A*** *. *. n6807 n3601
n3571 G******, T****** n6807 n3601
n4703 K****, S****-J****** n6807 n3601
n2146 K***, L**** n6807 n3601
Listing 2 represents the SPARQL query for the second competency question. In this case, the query selects DiagnosticMethod and PlasmaSource instances with the labels “Cavity Ring-Down Spectroscopy” and “kINPen-sci”, respectively, to infer the related datasets (of class Dataset). The Dataset class is connected with DiagnosticMethod and PlasmaSource classes mainly through PlasmaStudy. Therefore, the query first infers the plasma studies that are related to the selected DiagnosticMethod and PlasmaSource using the relevant properties, then finally, it infers the datasets related to the inferred plasma studies. Similarly, the answer of this query is shown in Table 2, which returns two datasets.
Listing 2: SPARQL query for “Which Dataset is obtained with DiagnosticMethod applied to diagnose PlasmaSource?”
PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>
PREFIX plasma-o: <http://plasma-mds.org/ontology/>
PREFIX vivo: <http://vivoweb.org/ontology/core#>
# Which datasets are obtained with the diagnostic
# method "Cavity Ring-Down Spectroscopy applied (CRDS)"
# to diagnose the plasma source "kINPen-sci?"
SELECT ?Dataset ?DatasetLabel ?PlasmaStudy ?DiagnosticMethod ?PlasmaSource
WHERE
# Define Dataset and PlasmaStudy
?Dataset rdf:type vivo:Dataset .
?PlasmaStudy rdf:type plasma-o:PlasmaStudy .
# Get the instance with label "Cavity Ring-Down Spectroscopy (CRDS)"
?DiagnosticMethod rdfs:label "Cavity Ring-Down Spectroscopy (CRDS)"@en-US .
# Get the instance with label "kINPen-sci"
?PlasmaSource rdfs:label "kINPen-sci"@en-US .
# Main query
?PlasmaStudy plasma-o:experimentUsesMethod ?DiagnosticMethod .
?PlasmaStudy plasma-o:hasPlasmaSource ?PlasmaSource .
?Dataset plasma-o:isOutputOf ?PlasmaStudy .
# Note: Dataset is connected to PlasmaStudy by the isOutputOf property,
# which is the inverse of the hasOutput property
# Showing the labels for Dataset
?Dataset rdfs:label ?DatasetLabel .
# Returns maximum of 5 results
LIMIT 5
Table 2: SPARQL results for “Which Dataset is obtained with DiagnosticMethod applied to diagnose PlasmaSource?” (Listing 2).
Dataset DatasetLabel PlasmaStudy DiagnosticMethod PlasmaSource
n5263 The spatial d… n4109 n6324 n6807
n1597 The localised… n1186 n6324 n6807
4 Conclusion and Outlook
The present work introduces the initial foundation of a domain-specific ontology for low-temperature plasma science and technology, named Plasma-O, and its application using the VIVO software to build a knowledge graph for the LTP field, named Plasma-KG. By aligning Plasma-O with the existing VIVO ontology, we expand VIVO’s capabilities to include specialized applications like the semantic cataloging of plasma sources and datasets, and the ability to address complex research questions through SPARQL queries. These enhanced capabilities directly contribute to the adoption of the FAIR data principles in the LTP field and connects the often disjunct topics of current research information management, research data management and semantic knowledge management in an interoperable way. The improved discoverability and findability of complex data facilitated by Plasma-O and VIVO significantly enhance the Findability and Accessibility aspects of FAIR data principles. Moreover, the standardized, machine-readable format of ontologies and knowledge graphs inherently promotes Interoperability and Reusability of LTP knowledge and research data. The presented methods provide a generalizable platform for domain-specific semantics, which can also be used by third-party services, e.g. to annotate research assets such as patents and scientific literature [Aras et al. (2024)]. Further, this work aims to initiate related discussions within the LTP community, hopefully steering the future development of the ontology and knowledge graph.
Currently, Plasma-O models general concepts in the LTP field, which limits the types and complexity of competency questions that can be formulated. In future work, we intend to expand the ontology to cover fine-grained details of plasma studies, such as properties of devices and samples. This could be relevant not only for knowledge discovery but also during the experimentation stage of research. Consequently, we plan to integrate the ontology and knowledge graph into an RDM tool like Adamant [^9] https://github.com/plasma-mds/adamant [Chaerony Siffa et al. (2022)]. This would allow researchers to interact with the ontology, thus the knowledge graph, as early as the experimentation stage (e.g. while documenting the experiments in the electronic laboratory notebook), enabling them to receive suggestions or reuse existing metadata for their experiments.
Code and Data Availability
The VIVO software version 1.13 is used in this work and publicly available at the corresponding GitHub release page [^10] https://github.com/vivo-project/VIVO/releases/tag/vivo-1.13.0. The Plasma Ontology and the example SPARQL queries presented in this work are publicly available at the corresponding GitHub page  [^11] https://github.com/plasma-mds/plasma-ontology. The plasma ontology (version 0.5.0) and knowledge graph used in this work are available at: https://doi.org/10.5281/zenodo.13325226. The VIVO software platform containing the plasma knowledge graph is currently under preparation and will be openly available following an embargo at: https://vivo.plasma-mds.org.
Acknowledgment
This work was partly funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—Project Number 496963457.